# Oligomorphic groups, categories of partial bijections, and ultrahomogeneous cubic spaces over finite fields

### v2

@@ 2. Oligomorphic groups and their trains
$4^{\lozenge}.$ ([38], [40]) Consider the group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ consisting of infinite invertible matrices over a finite field $\mathbb{F}$, which have finite number of elements in each row and each column (inverse matrices must have the same form). Let $V$ be as in the previous example, let $V^{\prime}$ be a copy of $V$. Let $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ act in $V\cup V^{\prime}$ by the formula $v\oplus v^{\prime}\mapsto gv\oplus g^{t-1}v^{\prime}$, where t denotes the transposition. Notice that this action preserves the natural pairing $V\times V^{\prime}\to\mathbb{F}$ defined by $\{v,v^{\prime}\}=\sum v_{j}v^{\prime}_{j}$. The group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ is a perfect oligomorphic subgroup in the symmetric group $S(V\cup V^{\prime})$, any perfect subset in $V\cup V^{\prime}$ is a union $X\cup X^{\prime}$ of a subspace $X\subset V$ and a subspace $X^{\prime}\subset V^{\prime}$. Morphisms $(X,X^{\prime})\to(Y,Y^{\prime})$ are pairs of partial linear bijections $\mu:X\to X^{\prime}$, $\nu:Y\to Y^{\prime}$ such that for any $v\in\operatorname{dom}\mu$, $v^{\prime}\in\operatorname{dom}\nu$ we have $\{\mu(v),\nu(v^{\prime})\}=\{v,v^{\prime}\}$.
@@
Consider a pair $(X,X^{\prime})$ of finite-dimensional linear spaces over $\mathbb{F}$ and a bilinear form $\lambda:X\times X^{\prime}\to\mathbb{F}$. We consider the subgroup $\operatorname{GL}(X\Join X^{\prime},\lambda)$ in $\operatorname{GL}(X)\times\operatorname{GL}(X^{\prime})$ preserving $\lambda$. Irreducible representations of $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ (and of its train) are enumerated by the following data. We choose a pair $(X,X^{\prime})$ with $\lambda$ and an irreducible representation of the group $\operatorname{GL}(X\Join X^{\prime},\lambda)$.
@@
$5^{\lozenge}.$ Let $\mathbb{F}$ be a finite field of characteristic $\neq 2$. Consider a non-degenerate quadratic form $\upsilon$ on $V$, all such forms are equivalent, so we can assume $\lambda(v)=\sum v_{j}^{2}$. Consider the group $\mathrm{O}(V)$ of linear transformations preserving this form. It is perfect oligomorphic on $V$, perfect subsets are finite-dimensional linear subspaces in $V$ (the proof for cubic forms from the next section is valid for quadratic forms). Objects of the train are finite-dimensional spaces $X$ equipped with quadratic forms. A morphisms $X\to Y$ is a bijective linear map from a subspace $A\subset X$ to a subspace $B\subset Y$ preserving the form.
@@
For the group $K=\operatorname{Aut}({\mathbb{Q}},<)$ double cosets $K(A)\backslash K/K(B)$ are enumerated by the following data. We consider an ordered set $Q$ and two order preserving injective maps $\lambda:A\to Q$, $\mu:B\to Q$ such that $\lambda(A)\cup\mu(B)=Q$. Taking the product $\lambda^{*}\mu$ in the sense of partial bijections we get the corresponding element of $[\![K(A)\backslash K/K(B)]\!]$.
@@
2) As we mentioned in Introduction, theory of oligomorphic groups arose not in representation theory (see [8], [40], [4]). From the point of view of refined representation theory (as semisimple Lie groups, symmetric groups, Chevalley groups, and infinite-dimensional groups) their unitary representations can seem exotic from one hand and relatively poor from the other hand.
We have seen a classifications of irreducible unitary representations of the group $\operatorname{Aut}(\mathcal{R})$, which includes classification of all irreducible representations of all finite groups. In this sense, the situation looks very complicated and very rich. Nevertheless, visible to the author problems of representation theory for the group $\operatorname{Aut}(\mathcal{R})$ are reduced using the Mackey machinery of induced representations [22], [23], [2] to problems for finite groups (which in this generality are unsolvable). Certainly, it is nice and impressive that the group $\operatorname{Aut}(\mathcal{R})$ admits a classification of unitary representations. But it is not clear how to continue this, if we want a continuation in the representation-theoretical spirit.
@@
However, Lieberman classification theorem (see [20], [35], [25]) for representations of $S_{\infty}$ is an important element of a wider ’representation theory of infinite symmetric groups’ (see, for example, [35], [19], [27], [7]). This theory was initiated by the work of Elmar Thoma [39] (1964) on ’characters of infinite symmetric group’. Olshanski [35] explained that the Thoma characters correspond to unitary representations of the following ’bisymmetric group’: we consider the subgroup in $S_{\infty}\times S_{\infty}$ consisting of pairs $(g_{1},g_{2})$ such that $g_{1}g_{2}^{-1}$ are finitely supported permutations [^10] Consider an irreducible unitary representation $\rho$ of a compact group $K$ in a Euclidean space $V$. Consider the space $\mathop{\mathrm{Op}}\nolimits(V)$ of all linear operators in $V$ equipped with the inner product $\langle A,B\rangle:=\mathop{\mathrm{tr}}\nolimits AB^{*}$. The group $K\times K$ acts in $\mathop{\mathrm{Op}}\nolimits(V)$ by operators $R(k_{1},k_{2})\,A=\rho(k_{1})A\rho(k_{2})^{-1}$. The identity operator $1_{V}$ is a unique vector fixed by the diagonal subgroup $\mathop{\mathrm{diag}}\nolimits(K)\subset K\times K$. The matrix element of $1_{V}$ is $\langle R(k_{1}k_{2})1_{V},1_{V}\rangle=\mathop{\mathrm{tr}}\nolimits\rho(k_{1}k_{2}^{-1})$. It is a value of the character of $\rho$ at $(k_{1}k_{2}^{-1})$. So (in the standard terminology) ’characters of a group $K$ are functions on $K\times K$ spherical with respect to $\mathop{\mathrm{diag}}\nolimits(K)$. It turns out that a pass to infinite limit in theories of characters (and, more generally, theories of spherical functions) is more interesting than attempts of straightforward generalizations of representations (’bisymmetric group’ arises in such limit). Also, it turns out that theory of Thoma characters can be included to a wider picture that (at least now) has no counterpart for finite symmetric groups, see [27].. This group contains the diagonal consisting of pairs $(g,g)$, where $g\in S_{\infty}$. There is a big zoo of other interesting groups containing $S_{\infty}$, see [27].
@@
4) Unitary representations of inverse limits $\lim\limits_{\leftarrow}G_{n}$ of oligomorphic groups are easily reduced to representation of prelimit oligomorphic groups $G_{n}$, see [40]. Precisely, any irreducible representation skips through one of groups $G_{n}$ (moreover, this is a general statement for any inverse limits of groups). But here we get new possibilities.
For example, consider the Bruhat–Tits tree $\mathcal{T}_{\infty}$ of infinite order, i.e., the simplicial tree whose vertices have countable valences. Fix a vertex $\xi$ and consider a ’circle’ $(\mathcal{T}_{\infty}^{n},\xi)\subset\mathcal{T}_{\infty}$ of radius $n$ with center at $\xi$. Consider the group $\operatorname{Aut}(\mathcal{T}_{\infty})$ of all automorphisms of $\mathcal{T}_{\infty}$, it subgroup $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$ fixing $\xi$, and the groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$. Groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$ are oligomorphic, the group $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$ is their inverse limit,
@@
An object of the corresponding category is a rooted tree $(L,\xi)$ such that distances between any vertex of $L$ and the root $\xi$ are $\leqslant n$. A morphism $(L,\xi)\to(M,\xi)$ is an isomorphism of a rooted subtree $(A,\xi)\subset(L,\xi)$ to a rooted subtree $(B,\xi)\subset(M,\xi)$.
We get a more rich situation, when we pass to natural overgroups of $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$, namely, to $\operatorname{Aut}(\mathcal{T}_{\infty})$, see [^11] The train of $\operatorname{Aut}(\mathcal{T}_{\infty})$ is not the category of finite trees and their partial isomorphisms as one might think looking to the previous examples, see [25], Sect VIII.6. [34], and to the group of spheromorphisms of $\mathcal{T}_{\infty}$, [28].
@@
Clearly, for groups (2.4) there are natural overgroups.
@@ 3. The group of automorphisms of the universal cubic spaces
3.3. Existence of the universal cubic space. Consider a linear space $Z$ over $\mathbb{F}$ of countable dimension, fix a basis $e_{1}$, $e_{2}$, …. For each non-ordered triple of vectors we assign a random value $\zeta(e_{i},e_{j},e_{k})\in\mathbb{F}$ uniformly distributed in $\mathbb{F}$. We assume that all these values are independent. We get a random cubic form $\zeta$ on $V$. We also can say that we consider a countable direct product of copies of the field $\mathbb{F}$ equipped with a uniform measure, factors are enumerated by non-ordered triples of basis elements $(e_{i},e_{j},e_{k})$, which can coincide.
@@
We claim that with probability 1 a form $\zeta$ satisfies the Urysohn extension property. Indeed, consider a finite dimensional cubic space $W$ and its subspace $V$ of codimension 1. Fix an isocubic embedding $\iota:V\to(Z,\zeta)$. Fix a basis $e_{1}$, …, $e_{k}$ in $\iota(V)$ and complete it to a basis $\{e_{\alpha}\}$ in $Z$. For each $\alpha>k$ we consider the subspace $W_{\alpha}$ generated by $\iota(V)$ and $e_{\alpha}$. The cubic form on $W_{\alpha}$ is determined by a collection of $k(k+1)/2+k+1$ numbers
@@
Now the universality is obvious – we take a basis $f_{1}$, …, $f_{k}$ in a cubic space $Y$, denote by $Y_{j}\subset Y$ the linear span of $f_{1}$, …, $f_{j}$. By the Urysohn property we can extend an isocubic embedding $Y_{j}\to\mathcal{Z}$ to an isocubic embedding $Y_{j+1}\to\mathcal{Z}$. We start with $Y_{0}=0$ and come to $Y_{k}=Y$.
Let us verify the ultrahomogeneity. Let $P_{k}$, $Q_{k}$ be two $k$-dimensional isocubic subspaces in a space $\mathcal{Z}$ satisfying the Urysohn property. Consider a sequence $a_{j}$ containing once all elements of $\mathcal{Z}\setminus P_{k}$ and a sequence $b_{j}$ containing once all elements of $\mathcal{Z}\setminus Q_{k}$.
@@
It remains to verify the uniqueness of $\mathcal{Z}$. We take two cubic spaces $\mathcal{Z}_{1}$, $\mathcal{Z}_{2}$ satisfying the Urysohn property and repeat the arguments from the proof of the ultrahomogeneity starting from $P_{0}=0$ in $\mathcal{Z}_{1}$ and $Q_{0}=0$ in $\mathcal{Z}_{2}$. $\square$
@@
Thus, the subgroup $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\subset S(\mathcal{Z})$ is perfect oligomorphic, perfect subsets are linear subspaces in $\mathcal{Z}$. So, Theorem 2.1 gives us a classification of unitary representations. By Theorem 2.5, we have a well-defined category of reduced double cosets, it is equivalent to the category of partial linear isocubic bijections.
@@
Objects of the category $\operatorname{Tra}\bigl{(}\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\bigr{)}$ are finite-dimensional cubic linear spaces. A morphism $V\to W$ is an isocubic linear bijection of a linear subspace $A\subset V$ to a linear subspace $B\subset W$.
@@
is direct and has countable codimension in $\mathcal{Z}$
@@
In the notation of Lemma 3.8, denote $W:=V\cap V^{\prime}$. Decompose $V=W\dotplus X$, $V^{\prime}=W\dotplus Y$. Let $W\dotplus\widetilde{X}\dotplus\widetilde{Y}$ be a subspace in $\mathcal{Z}$ such that there is an isocubic map $r:W\oplus X\oplus Y\to W\oplus\widetilde{X}\oplus\widetilde{Y}$ identical on $W$ and sending $X\to\widetilde{X}$, $Y\to\widetilde{Y}$. Then there is an element $h\in H$ such that the restriction of $h$ to $W\oplus X\oplus Y$ is $r$.
Proof of Lemma 3.9. We have $K(W\oplus X)$, $K(W\oplus Y)\subset H$. Notice that
@@
Step 2. We once again apply Lemma 3.7 and keep in mind its proof. Choose bases
@@
Let $j$ range in $1$, $2$, $3$, …, and $w_{\alpha}$, $x_{\beta}$, $y_{\gamma}$ range in sets of basis vectors in $W$, $X$, $Y$. Then the following collection of $\mathbb{F}$-valued random variables
@@
$\sigma:W\oplus X\oplus Y^{j_{*}}\to W\oplus rX\oplus Y^{j_{*}},$
which is identical on $W\oplus Y^{j_{*}}$ and equals $r$ on $X$ (and therefore $\sigma$ equals $r$ on $W\oplus X$). This map is $\zeta$-isocubic. Indeed,
1) by definition, it is isocubic on $W\oplus Y^{j_{*}}$;
2) $r$ is isocubic on $W\oplus X$;
3) by (3.6)-(3.7) it is isocubic on $X\oplus Y^{j_{*}}$;
@@
The map $\sigma$ extends to an element $\widetilde{\sigma}$ of the subgroup $K(W\oplus Y^{j_{*}})\subset H$. Conjugating $K(W\oplus X)$ by $\widetilde{\sigma}$, we get that $K(W\oplus\widetilde{X})\subset H$. Also, by the construction, we have the following isocubic maps
@@
So, we have an isocubic map $W\oplus\widetilde{X}\oplus Y^{j_{*}}\to W\oplus\widetilde{X}\oplus\widetilde{Y}$. By the construction, it fixes $W\oplus\widetilde{X}$. Therefore there is an element $\xi$ of $K(W\oplus\widetilde{X})$ sending one subspace to another. We take $h=\xi\widetilde{\sigma}$square

### v3

@@ 2. Oligomorphic groups and their trains
$4^{\lozenge}.$ ([39], [41]) Consider the group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ consisting of infinite invertible matrices over a finite field $\mathbb{F}$, which have finite number of elements in each row and each column (inverse matrices must have the same form). Let $\ell(\mathbb{F})$ be as in the previous example, let $\ell^{\prime}(\mathbb{F})$ be a copy of $V$. Let $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ act in $\ell(\mathbb{F})\sqcup\ell^{\prime}(\mathbb{F})$ by the formula $v\mapsto gv$, $v^{\prime}\mapsto g^{t-1}v^{\prime}$, where $v\in\ell(\mathbb{F})$, $v^{\prime}\in\ell^{\prime}(\mathbb{F})$, and t denotes the transposition. Notice that this action preserves the natural pairing $\ell(\mathbb{F})\times\ell^{\prime}(\mathbb{F})\to\mathbb{F}$ defined by $\{v,v^{\prime}\}=\sum v_{j}v^{\prime}_{j}$. The group $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ is a perfect oligomorphic subgroup in the symmetric group $S(\ell(\mathbb{F})\sqcup\ell^{\prime}(\mathbb{F}))$, any perfect subset in $\ell(\mathbb{F})\sqcup\ell^{\prime}(\mathbb{F})$ is a union $X\sqcup X^{\prime}$ of a subspace $X\subset\ell(\mathbb{F})$ and a subspace $X^{\prime}\subset\ell^{\prime}(\mathbb{F})$. Morphisms $(X,X^{\prime})\to(Y,Y^{\prime})$ are pairs of partial linear bijections $\mu:X\to Y^{\prime}$, $\nu:X^{\prime}\to Y^{\prime}$ such that for any $v\in\operatorname{dom}\mu$, $v^{\prime}\in\operatorname{dom}\nu$ we have $\{\mu(v),\nu(v^{\prime})\}=\{v,v^{\prime}\}$.
We can consider the following equivalent category. Objects are pairs $(X,X^{\prime})$ of finite-dimensional linear spaces over $\mathbb{F}$ equipped with a bilinear form $\Phi:X\times X^{\prime}\to\mathbb{F}$. Morphisms $(X,X^{\prime};\Phi)\to(Y,Y^{\prime};\Psi)$ are pairs of linear partial bijections $\lambda:X\to Y$ and $\mu:X^{\prime}\to Y^{\prime}$ such that $\Psi(\lambda(x),\mu(x^{\prime}))=\Phi(x,x^{\prime})$.
@@
We consider the subgroup $\operatorname{GL}(X\Join X^{\prime},\Phi)$ in $\operatorname{GL}(X)\times\operatorname{GL}(X^{\prime})$ preserving $\Phi$. Irreducible representations of $\operatorname{GL}^{\diamond}(\infty,\mathbb{F})$ (and of its train) are enumerated by the following data (up to the natural equivalence): a pair $(X,X^{\prime})$ equipped with $\Phi$ and an irreducible representation of the group $\operatorname{GL}(X\Join X^{\prime},\Phi)$.
Notice that pairs $(X,X^{\prime};\Phi)$ are classified (up to a natural equivalence) by $\xi=\dim X$, $\xi^{\prime}=\dim X^{\prime}$, and the rank $\rho$ of $\Phi$. The group $\operatorname{GL}(X\Join X^{\prime},\Phi)$ is a semidirect product
$\Bigl{(}\operatorname{GL}(\xi-\rho,\mathbb{F})\times\operatorname{GL}(\xi^{\prime}-\rho,\mathbb{F})\times\operatorname{GL}(\rho,\mathbb{F})\Bigr{)}\ltimes\Bigl{(}(\mathbb{F}^{\rho}\otimes\mathbb{F}^{\xi-\rho})\times(\mathbb{F}^{\xi^{\prime}-\rho}\otimes\mathbb{F}^{\rho})\Bigr{)}.$
For a classification of irreducible representations of $\operatorname{GL}(X\Join X^{\prime},\Phi)$ we can apply Mackey’s arguments [23] but they lead to a branching recursive problems and do not give a final answer.
@@
$5^{\lozenge}.$ Let $\mathbb{F}$ be a finite field of characteristic $\neq 2$. Consider a non-degenerate quadratic form $\Phi$ on $\ell(\mathbb{F})$, all such forms are equivalent [^11] Let $\varepsilon\in\mathbb{F}$ be a quadratic non-residue. We reduce a quadratic form to a sum of squares $\sum a_{j}w_{j}^{2}$, where $a_{j}$=1 or $\varepsilon$, in the usual way. Binary forms $z_{1}^{2}+z_{2}^{2}$ and $\varepsilon z_{1}^{2}+\varepsilon z_{2}^{2}$ are equivalent. Therefore, each form is equivalent to the form $\sum w_{2j}^{2}+\sum\varepsilon w_{2j+1}^{2}$., so we can assume $\Phi(v)=\sum v_{j}^{2}$. Consider the group $\mathrm{O}(\infty)$ of linear transformations preserving this form. It is perfect oligomorphic on $\ell(\mathbb{F})$, perfect subsets are finite-dimensional linear subspaces in $\ell(\mathbb{F})$ (the proof for cubic forms from the next section is valid for quadratic forms). A morphism $X\to Y$ is a bijective linear map $\nu$ from a subspace $A\subset X$ to a subspace $B\subset Y$ such that $\Phi(\nu(x_{1}),\nu(x_{2}))=\Phi(x_{1},x_{2})$.
We can also consider the following equivalent category. Its objects are finite dimensional linear spaces $V$ over $\mathbb{F}$ equipped with quadratic forms $\Phi_{V}$, morphisms are partial linear bijections preserving forms. Irreducible unitary representation are enumerated by quadratic spaces $(V,\Phi_{V})$ defined up to the equivalence (i.e., by $\dim V$, rank of $\Phi_{V}$, and the discriminant of $\Phi_{V}$) and by irreducible representations of groups $\rm{O}(V,\Phi_{V})$ of isometries of $(V,\Phi_{V})$.
@@
For the group $K=\operatorname{Aut}({\mathbb{Q}},<)$, double cosets $K(A)\backslash K/K(B)$ are enumerated by the following data. We consider an ordered set $Q$ and two order preserving injective maps $\lambda:A\to Q$, $\mu:B\to Q$ such that $\lambda(A)\cup\mu(B)=Q$. Taking the product $\lambda^{*}\mu$ in the sense of partial bijections we get the corresponding element of the reduced set $[\![K(A)\backslash K/K(B)]\!]$.
@@
2) As we mentioned in Introduction, theory of oligomorphic groups arose not in representation theory (see [8], [41], [4]). From the point of view of refined representation theory (as semisimple Lie groups, symmetric groups, Chevalley groups, $p$-adic groups, and infinite-dimensional groups) their unitary representations can seem exotic from one hand and relatively poor from the other hand.
We have seen a classifications of irreducible unitary representations of the group $\operatorname{Aut}(\mathcal{R})$, which includes classification of all irreducible representations of all finite groups. In this sense, the situation looks very complicated and very rich. Nevertheless, visible to the author problems of representation theory for the group $\operatorname{Aut}(\mathcal{R})$ are reduced using the Mackey machinery of induced representations [22], [24], [2] to problems for finite groups (which in this generality are unsolvable). Certainly, it is nice and impressive that the group $\operatorname{Aut}(\mathcal{R})$ admits a classification of unitary representations. But it is not clear how to continue this, if we want a continuation in the representation-theoretical spirit (but see [42] on actions on measure spaces and de Finetti type theorems).
@@
However, Lieberman classification theorem (see [20], [36], [26]) for representations of $S_{\infty}$ is an important element of a wider ’representation theory of infinite symmetric groups’ (see, for example, [36], [19], [28], [7]). This theory was initiated by the work of Elmar Thoma [40] (1964) on ’characters of infinite symmetric group’. Olshanski [36] explained that the Thoma characters correspond to unitary representations of the following ’bisymmetric group’: we consider the subgroup in $S_{\infty}\times S_{\infty}$ consisting of pairs $(g_{1},g_{2})$ such that $g_{1}g_{2}^{-1}$ are finitely supported permutations [^12] Consider an irreducible unitary representation $\rho$ of a compact group $K$ in a Euclidean space $V$. Consider the space $\mathop{\mathrm{Op}}\nolimits(V)$ of all linear operators in $V$ equipped with the inner product $\langle A,B\rangle:=\mathop{\mathrm{tr}}\nolimits AB^{*}$. The group $K\times K$ acts in $\mathop{\mathrm{Op}}\nolimits(V)$ by operators $R(k_{1},k_{2})\,A=\rho(k_{1})A\rho(k_{2})^{-1}$. The identity operator $1_{V}$ is a unique vector fixed by the diagonal subgroup $\mathop{\mathrm{diag}}\nolimits(K)\subset K\times K$. The matrix element of $1_{V}$ is $\langle R(k_{1}k_{2})1_{V},1_{V}\rangle=\mathop{\mathrm{tr}}\nolimits\rho(k_{1}k_{2}^{-1})$. It is a value of the character of $\rho$ at $(k_{1}k_{2}^{-1})$. So (in the standard terminology) ’characters of a group $K$ are functions on $K\times K$ spherical with respect to $\mathop{\mathrm{diag}}\nolimits(K)$. It turns out that a pass to infinite limit in theories of characters (and, more generally, in theories of spherical functions) is more interesting than attempts of straightforward generalizations of representations (’bisymmetric group’ arises in such limit). Also, it turns out that theory of Thoma characters can be included to a wider picture that (at least now) has no counterpart for finite symmetric groups, see [28].. This group contains the diagonal consisting of pairs $(g,g)$, where $g\in S_{\infty}$. There is a big zoo of other interesting groups containing $S_{\infty}$, see [28].
@@
4) Unitary representations of inverse limits $\lim\limits_{\leftarrow}G_{n}$ of oligomorphic groups are easily reduced to representation of prelimit oligomorphic groups $G_{n}$, see [41]. Precisely, any irreducible representation skips through one of groups $G_{n}$ (moreover, this is a general statement for any inverse limits of groups, this follows from [26], Proposition VIII.1.3). But here we get new possibilities.
For example, consider the Bruhat–Tits tree $\mathcal{T}_{\infty}$ of infinite order, i.e., the simplicial tree whose vertices have countable valences. Fix a vertex $\xi$ and consider a ’ball’ $(\mathcal{T}_{\infty}^{n},\xi)\subset\mathcal{T}_{\infty}$ of radius $n$ with center at $\xi$. Consider the group $\operatorname{Aut}(\mathcal{T}_{\infty})$ of all automorphisms of $\mathcal{T}_{\infty}$, its subgroup $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$ fixing $\xi$, and the groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$. The groups $\operatorname{Aut}(\mathcal{T}_{\infty}^{n},\xi)$ are oligomorphic, the group $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$ is their inverse limit,
@@
An object of the corresponding category is a finite rooted tree $(L,\xi)$ such that distances between any vertex of $L$ and the root $\xi$ are $\leqslant n$. A morphism $(L,\xi)\to(M,\xi)$ is an isomorphism of a rooted subtree $(A,\xi)\subset(L,\xi)$ to a rooted subtree $(B,\xi)\subset(M,\xi)$.
We get a more rich situation, when we pass to the natural overgroup of $\operatorname{Aut}(\mathcal{T}_{\infty},\xi)$, namely, to $\operatorname{Aut}(\mathcal{T}_{\infty})$, see [^13] The train of $\operatorname{Aut}(\mathcal{T}_{\infty})$ is not the category of finite trees and their partial isomorphisms as one might think looking to the previous examples, see [26], Sect VIII.6. [35], and to the group of spheromorphisms of $\mathcal{T}_{\infty}$, [29].
@@
Clearly, the groups (2.4) have natural overgroups consisting of $p$-adic matrices of the same structure.
@@ 3. The group of automorphisms of the universal cubic spaces
3.3. Existence of the universal cubic space. A. Consider a linear space $Z$ over $\mathbb{F}$ of countable dimension, fix a basis $e_{1}$, $e_{2}$, …. For each non-ordered triple of vectors we assign a random value $\zeta(e_{i},e_{j},e_{k})\in\mathbb{F}$ uniformly distributed in $\mathbb{F}$. We assume that all these values are independent. We get a random cubic form $\zeta$ on $V$. We also can say that we consider a countable direct product of copies of the field $\mathbb{F}$ equipped with a uniform measure, factors are enumerated by non-ordered triples of basis elements $(e_{i},e_{j},e_{k})$, which can coincide.
@@
B. We claim that with probability 1 a form $\zeta$ satisfies the Urysohn extension property. Indeed, consider a finite dimensional cubic space $W$ and its subspace $V$ of codimension 1. Fix an isocubic embedding $\iota:V\to(Z,\zeta)$. Fix a basis $e_{1}$, …, $e_{k}$ in $\iota(V)$ and complete it to a basis $\{e_{\alpha}\}$ in $Z$. For each $\alpha>k$ we consider the subspace $W_{\alpha}$ generated by $\iota(V)$ and $e_{\alpha}$. The cubic form on $W_{\alpha}$ is determined by a collection of $k(k+1)/2+k+1$ numbers
@@
C. Now the universality is obvious – we take a basis $f_{1}$, …, $f_{k}$ in a cubic space $Y$, denote by $Y_{j}\subset Y$ the linear span of $f_{1}$, …, $f_{j}$. By the Urysohn property we can extend an isocubic embedding $Y_{j}\to\mathcal{Z}$ to an isocubic embedding $Y_{j+1}\to\mathcal{Z}$. We start with $Y_{0}=0$ and come to $Y_{k}=Y$.
D. Let us verify the ultrahomogeneity. Let $P_{k}$, $Q_{k}$ be two $k$-dimensional isocubic subspaces in a space $\mathcal{Z}$ satisfying the Urysohn property. Consider a sequence $a_{j}$ containing once all elements of $\mathcal{Z}\setminus P_{k}$ and a sequence $b_{j}$ containing once all elements of $\mathcal{Z}\setminus Q_{k}$.
@@
E. It remains to verify the uniqueness of $\mathcal{Z}$. We take two cubic spaces $\mathcal{Z}_{1}$, $\mathcal{Z}_{2}$ satisfying the Urysohn property and repeat the arguments from the proof of the ultrahomogeneity starting from $P_{0}=0$ in $\mathcal{Z}_{1}$ and $Q_{0}=0$ in $\mathcal{Z}_{2}$. $\square$
@@
Thus, the subgroup $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\subset S(\mathcal{Z})$ is perfect oligomorphic, perfect subsets are linear subspaces in $\mathcal{Z}$. So, Theorem 2.1 gives us a classification of unitary representations of $\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})$. By Theorem 2.5, we have a well-defined category of reduced double cosets, it is equivalent to the category of partial linear isocubic bijections.
@@
The category $\operatorname{Tra}\bigl{(}\mathop{\mathrm{Isoc}}\nolimits(\mathcal{Z})\bigr{)}$ is equivalent to the following category: its objects are finite-dimensional cubic linear spaces. A morphism $V\to W$ is an isocubic linear bijection of a linear subspace $A\subset V$ to a linear subspace $B\subset W$.
@@
is direct and has countable codimension in $\mathcal{Z}$.
@@
In the notation of Lemma 3.8, denote $W:=V\cap V^{\prime}$. Decompose $V=W\dotplus X$, $V^{\prime}=W\dotplus Y$. Let $W\dotplus\widetilde{X}\dotplus\widetilde{Y}$ be a subspace in $\mathcal{Z}$ such that there exists an isocubic map $r:W\dotplus X\dotplus Y\to W\dotplus\widetilde{X}\dotplus\widetilde{Y}$ identical on $W$ and sending $X\to\widetilde{X}$, $Y\to\widetilde{Y}$. Then there is an element $h\in H$ such that the restriction of $h$ to $W\dotplus X\dotplus Y$ is $r$.
Proof of Lemma 3.9. We have $K(W\dotplus X)$, $K(W\dotplus Y)\subset H$. Notice that
@@
Step 2. So, we assume that (3.4) holds. Now we intend to apply Lemma 3.7 and Corollary 3.3 again. Choose bases
@@
Let $j$ range in $1$, $2$, $3$, …, and $w_{\alpha}$, $x_{\beta}$, $y_{\gamma}$ range in sets of basis vectors of $W$, $X$, $Y$ respectively. Then the following collection of $\mathbb{F}$-valued random variables
@@
$\sigma:W\dotplus X\dotplus Y^{j_{*}}\to W\dotplus rX\dotplus Y^{j_{*}},$
which is identical on $W\dotplus Y^{j_{*}}$ and equals $r$ on $X$ (and therefore $\sigma$ equals $r$ on $W\dotplus X$). This map is $\zeta$-isocubic. Indeed,
1) by definition, it is isocubic on $W\dotplus Y^{j_{*}}$;
2) $r$ is isocubic on $W\dotplus X$;
3) by (3.6)-(3.7) it is isocubic on $X\dotplus Y^{j_{*}}$;
@@
The map $\sigma$ extends to an element $\widetilde{\sigma}$ of the subgroup $K(W\dotplus Y^{j_{*}})\subset H$. Conjugating $K(W\dotplus X)$ by $\widetilde{\sigma}$, we get that $K(W\dotplus\widetilde{X})\subset H$. Also, by the construction, we have the following isocubic maps
@@
So, we have an isocubic map $W\dotplus\widetilde{X}\dotplus Y^{j_{*}}\to W\dotplus\widetilde{X}\dotplus\widetilde{Y}$. By the construction, it fixes $W\dotplus\widetilde{X}$. Therefore there is an element $\xi$ of $K(W\dotplus\widetilde{X})$ sending one subspace to another. We take $h=\xi\widetilde{\sigma}$square